OpenStudy (anonymous):

what has an x-intercept of 6 and a y-intercept of 3

OpenStudy (anonymous):

are you supposed to find an equation?

OpenStudy (anonymous):

i think so i need the answer asap

OpenStudy (anonymous):

do you understand what slope intercept form is?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

have you learned it yet?

OpenStudy (anonymous):

ya i just forgot its been about two years scince i used it

OpenStudy (anonymous):

ooo okies

OpenStudy (anonymous):

is that a question

OpenStudy (anonymous):

yap

OpenStudy (anonymous):

do you understan what all those are x intercept y intercept slope

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

i just need the answer asap

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

so it all begins with slope you want slope you get y - y m = ---- x - x

OpenStudy (anonymous):

y is that?

OpenStudy (anonymous):

Y= Mx+b Y= 2x+3 I think

OpenStudy (anonymous):

sorry convert to slope intercept form y - y = m ( x - x ) m = slope y = y value x = x intercept y - y = m ( x - x ) + b b = y intercept